OpenStudy (anonymous):

how many gallons of water would it take to fill up 9.5 deep 15 foot wide 15 foot long

OpenStudy (anonymous):

first figure out the volume of the pool

OpenStudy (anonymous):

do you know the formula for a rectangular volume?

OpenStudy (anonymous):

\[V=LWH\] where L=length,W=Width, H=height

OpenStudy (anonymous):

so its 10times15times15?

OpenStudy (anonymous):

I got 2250

OpenStudy (anonymous):

9.5*15*15=2137.5

OpenStudy (anonymous):

oh oops ok now what

OpenStudy (anonymous):

since you multiplied foot*foot*foot, you're answer ( 2137.5 ) should be in cubic ft

OpenStudy (anonymous):

how do I get that in cubic feet

OpenStudy (anonymous):

it is in cubic feet

OpenStudy (anonymous):

oh im stupid

OpenStudy (anonymous):

\[V=15ft*15ft*9.5=2137.5(ft)(ft)(ft)=2137.5ft^3\]

OpenStudy (anonymous):

next use dimensional analysis do you know what this is?

OpenStudy (anonymous):

noo sorry

OpenStudy (anonymous):

alright basically dimensional analysis is using proportions/ratios of measurement to switch to different types of measurement

OpenStudy (anonymous):

i'll give you an example, say you have 5 inches and you want to get this into centimeters you know that 1inch = 2.14 cm so the ratio can be written either as \[\frac{1inch}{2.14cm}\] or \[\frac{2.14cm}{1inch}\]

OpenStudy (anonymous):

these ratios will be used to cancel out terms first start out with whta they give you, my example it is 5 inches